# A Non-Commutative Kalb-Ramond Black Hole

### v2

@@ IV.2 Fermionic particle modes
This section investigates the quantum tunneling of fermions across the event horizon associated with the black hole geometry under consideration. While various approaches—such as those using generalized Painlevé–Gullstrand or Kruskal–Szekeres coordinates—have been explored in earlier foundational studies o69 , our analysis starts from the general metric form provided in Eq. (24). In a curved spacetime background, the evolution of spin–$\tfrac{1}{2}$ particles is governed by the Dirac equation, which takes the following form:
@@
$\gamma^{\theta}(\Theta,\ell)$ $=$ $\frac{1}{\sqrt{{C^{(\Theta,\ell)}}}}\left(\begin{array}[]{cc}\bf{0}&\sigma_{1}\\ \sigma_{1}&\bf{0}\end{array}\right),\;\;\gamma^{\varphi}(\Theta,\ell)=\frac{1}{{D^{(\Theta,\ell)}}}\left(\begin{array}[]{cc}\bf{0}&\sigma_{2}\\ \sigma_{2}&\bf{0}\end{array}\right).$
@@
To represent a Dirac field whose spin is oriented upward along the positive $r$–direction, the following ansatz is employed vanzo2011tunnelling :
@@
The analysis centers on the spin–up (+)(+)( + ) configuration, corresponding to alignment along the positive $r$–axis. The spin-down ()(-)( - ) case, which represents orientation in the opposite direction, can be handled through an analogous procedure. Inserting the spinor ansatz from Eq. (58) into the Dirac equation yields the following set of expressions, in line with the method outlined by Vanzo et al. vanzo2011tunnelling
@@
$\begin{split}n_{\psi}(\Theta,\ell)&={\frac{\gamma_{\psi}(\Theta,\ell)}{1+\gamma_{\psi}(\Theta,\ell)}}\\ &=\,\,\frac{1}{\exp\left(\frac{128\pi(\ell-1)^{5}M^{3}\omega}{16(\ell-1)^{3}M^{2}-\Theta^{2}\left(9\ell^{2}+5\right)}\right)+1}.\end{split}$ (74)
Fig. 7 displays the fermionic particle creation density, $n_{\psi}$. As observed in the bosonic case, both the non–commutative parameter $\Theta$ and the Lorentz–violating parameter $\ell$ enhance the overall particle production rate, leading to a noticeable increase in the density’s amplitude therefore.
@@ V Evaporation and Emission rate
$\frac{\mathrm{d}M}{\mathrm{d}t}=-\alpha a\sigma T^{4},$ (75)
with $\sigma$ corresponds to the effective radiating (cross–section) area, $a$ is the radiation constant, $\alpha$ accounts for the greybody factor, and $T$ denotes the Hawking temperature. It is important to highlight that the emitted radiation predominantly consists of massless particles, such as photons and neutrinos hiscock1990evolution ; page1976particle . The cross–sectional area $\sigma$ is approximated by $\pi R_{sh}^{2}$, where $R_{sh}$ denotes the shadow radius, serving as the effective emission radius in the high–frequency (geometric optics) limit—often referred to as $\sigma_{\text{lim}}$. Under these conditions, the greybody factor simplifies to $\alpha\to 1$, as noted in liang2025einstein . Moreover, after some algebraic manipulations, we can write $\sigma_{lim}$ as follows:
[…]
@@
$\begin{split}\frac{\mathrm{d}M}{\mathrm{d}t}=&-\frac{\xi\left(3\Theta^{2}+16(\ell-1)^{2}(\ell+1)M^{2}\right)^{4}\left(\Theta^{2}(\ell+2)+72(3\ell-2)M^{2}\right)^{2}}{206158430208\pi^{3}(\ell-1)^{12}M^{14}}\\ \approx&\,\,-\frac{27\xi}{4096\left(\pi^{3}M^{2}\right)}-\frac{135\xi\ell}{4096\left(\pi^{3}M^{2}\right)}+\left(-\frac{465\xi\ell}{16384\pi^{3}M^{4}}-\frac{39\xi}{8192\left(\pi^{3}M^{4}\right)}\right)\Theta^{2}\,,\end{split}$ (77)
where we expanded up to second order in $\Theta$ and first order in $\ell$, with $\xi=a\alpha$, the expression simplifies accordingly. The next step involves evaluating the following integral:
@@
Figure 9: The evaporation lifetime at the its final stage $M_{f}\to 0$ for different values of initial mass $M_{i}$ by considering different value of $\Theta$ (up panel) and different values of $\ell$ (bottom panel).
@@
Quantum effects manifest prominently in the vicinity of a black hole’s event horizon, where transient particle–antiparticle pairs emerge due to vacuum fluctuations. Occasionally, one of these particles—carrying positive energy—may traverse the potential barrier and escape via quantum tunneling. This escape mechanism leads to a net energy loss from the black hole, contributing to its gradual evaporation, a process commonly known as Hawking radiation. For an observer located far from the black hole, this evaporation is indirectly associated with the absorption profile at high energies. In such regimes, the absorption cross–section tends toward a limiting value, denoted by $\sigma_{\text{lim}}$. As reported in Refs. decanini2011universality ; papnoi2022rotating , this behavior governs the spectral energy emission, which can be formulated as follows:
$\frac{{{\mathrm{d}^{2}}E}}{{\mathrm{d}\omega\mathrm{d}t}}=\frac{{2{\pi^{2}}\sigma_{lim}}}{{{e^{\frac{\omega}{T^{(\Theta,\ell)}}}}-1}}{\omega^{3}},$ (81)
@@
$\frac{\mathrm{d}^{2}N}{\mathrm{d}\omega\mathrm{d}t}=\frac{2\pi^{2}\,\sigma_{lim}\,\omega^{2}}{e^{\,\omega/T^{(\Theta,\ell)}}-1}=\frac{\pi^{3}\omega^{2}\left(\Theta^{2}(\ell+2)+72(3\ell-2)M^{2}\right)^{2}}{384M^{2}\left(e^{-\frac{128\pi(\ell-1)^{3}M^{3}\omega}{3\Theta^{2}+16(\ell-1)^{2}(\ell+1)M^{2}}}-1\right)}.$ (83)
@@ VI Massless Scalar Field in a Non-Commutative Background
Substituting the Eq. (103)-(104) into the Eq. (102) and considering the new redefined radial function $\Psi_{lm}(r^{*})$ in the Klein–Gordon equation, a one–dimensional Schrödinger–like wave equation will be obtained as
@@
$\left.\left.+768r^{3}\left(\frac{f^{4}(r)}{2\sqrt{f(r)}}-\frac{3f^{\prime\prime}(r)^{2}}{4f(r)^{3/2}}-\frac{15f^{\prime}(r)^{4}}{16f(r)^{7/2}}-\frac{f^{3}(r)f^{\prime}(r)}{f(r)^{3/2}}+\frac{9f^{\prime}(r)^{2}f^{\prime\prime}(r)}{4f(r)^{5/2}}\right)\right)\right)\Big{]}.$ (108)
@@
where the primes denote derivatives with respect to the radial coordinate $r$. It is evident that setting $\Theta\to 0$ recovers the conventional effective potential for a scalar field in the Kalb–Ramond black hole background to $l=1$, as we should expect.
@@ VII Quasinormal modes
[…]
@@ VIII Time–domain solution
Here, the evolution begins by assigning the scalar field a vanishing value along the initial surface $v=v_{0}$, that is, $\psi(u,v_{0})=0$. This choice simplifies implementation without restricting the general applicability of the method. The numerical integration proceeds by holding $u$ constant while incrementally advancing in the $v$ direction, making use of the null boundary data. For the sake of simplicity, the scalar perturbation is treated in a background where $M=1$. The initial waveform is introduced as a Gaussian pulse centered at $v=0$, with a fixed width $\sigma=1$ and zero amplitude at the origin. The computational grid is defined over the square domain $[0,1000]\times[0,1000]$ in the $(u,v)$ plane, using a uniform spacing of $h=0.1$.
The progression of the scalar field $\psi$ over time is illustrated in Fig. 21, where distinct values of the non–commutative parameter $\Theta$—ranging from $0.1$ to $0.4$—are examined under the constraint $\ell=0.1$. The behavior is evaluated across different angular modes: $l=1$ (upper left), $l=2$ (upper right), and $l=3$ (bottom), with the azimuthal number fixed at $m=+1$. The evolution profiles indicate that the field tends to stabilize as time increases, eventually flattening out to small values.
@@
Figure 21: The time–dependent behavior of the field $\psi$ is illustrated for different angular indices, with panels corresponding to $l=1$ (top left), $l=2$ (top right), and $l=3$ (bottom), under the fixed choice of $m=+1$. The parameter $\Theta$ is varied across the set ${0.1,0.2,0.3,0.4}$ to examine its influence on the wave evolution, while the value of $\ell$ is kept unchanged at $0.1$ throughout the analysis.
@@ IX Estimated constraints from the solar system tests
Moreover, the expressions for the conserved quantities associated with the system—namely, the energy $E$ and the angular momentum $L$—can be written as follows
@@ IX.1 Perihelion precession of Mercury
$u_{0}=1+e\cos(\varphi)\,.$ (121)
Here, by inserting the expanded form of $u$ into Eq. (120), one arrives at the following differential equation structure:
@@
When considering the correction associated with the $\epsilon$ parameter in Eq. (122), its influence must vanish to preserve the perturbative structure, yielding new corrections. Among the various sinusoidal terms that arise—most of which average to zero over a full orbit—only one term, $-\frac{1}{2}e\varphi\sin(\varphi)$, contributes meaningfully by producing a secular variation. After accounting for the Newtonian background, the relativistic correction, and this new contribution, and by reintroducing the original theoretical parameters, the resulting expression for $u(\varphi)$ becomes:
@@
The angular momentum $L$ of a planetary orbit can be expressed in terms of the semi-major axis $a$ and eccentricity $e$ through the relation $L^{2}=Ma(1-e^{2})$, while the specific orbital energy per unit mass is given by $E=-M/(2a)$ Goldstein:2002 . In natural units, the physical parameters for Mercury’s orbit and the Sun are set as follows: the solar mass is $M=M_{\odot}=9.138\times 10^{37}$, the semi–major axis of Mercury’s orbit is $a=3.583\times 10^{45}$, and its eccentricity is $e=0.2056$. Using these values, one obtains $L=5.600\times 10^{41}$, confirming the validity of a perturbative approach due to the smallness of $M^{2}/L^{2}$ contributions. The squared energy, calculated as $E^{2}=1.627\times 10^{-16}$, is small enough that energy–dependent corrections can safely be disregarded.
With the parameter $\epsilon$ defined by $\epsilon=-\ell-\frac{\Theta^{2}}{2L^{2}}(1-E^{2})$, one can derive independent constraints on $\ell$ and $\Theta$. Setting $\Theta=0$ isolates the Kalb–Ramond contribution, leading to the constraint $-1.817\times 10^{-11}\leq\ell\leq 3.634\times 10^{-12}$. Alternatively, taking $\ell=0$ allows for bounding the non–commutative parameter, resulting in the range $-2976.57,\text{m}^{2}\leq\Theta^{2}\leq 595.315,\text{m}^{2}$. For completeness, both positive and negative branches of $\Theta^{2}$ are retained in the analysis.
@@ IX.2 Deflection of light
In this context, the ratio $L/E$ corresponds to the impact parameter $b$. To isolate the effects ascribed to $\Theta$, we first set $\ell=0$. Notably, the resulting corrections involving the parameter $\Theta$ are universal—they appear independently of the central mass $M$—and thus modify the Newtonian trajectory from the outset. By appropriately redefining the mass term, the leading–order contributions can be expressed as follows:
@@
General Relativity predicts a deflection angle of $\delta_{\text{GR}}=4M/b=1.7516687^{\prime\prime}$. Observationally, the measured deflection is expressed as $\delta_{\text{Exp}}=\frac{1}{2}(1+\gamma)\times 1.7516687^{\prime\prime}$, with $\gamma=0.99992\pm 0.00012$ dsasdas . To examine the influence of $\Theta$, the dimensionless factor $1-\Theta^{2}/(4b^{2})$ is compared with the empirical ratio $(1+\gamma)/2$. This comparison yields the following constraint: $-3.872\times 10^{13}\,\text{m}^{2}\leq\Theta^{2}\leq 1.936\times 10^{14}\,\text{m}^{2}$.
@@
By assuming the regime of small angular deviations, $\varphi\ll 1$, and following the same reasoning outlined previously, the resulting deflection angle in the presence of the Kalb–Ramond background becomes $\delta_{\ell}=\frac{4M}{b}(1-\frac{3}{2}\ell)$. This expression allows us to infer the constraint on $\ell$, yielding the range $-1.333\times 10^{-5}\leq\ell\leq 6.667\times 10^{-5}$.

### v3

@@ IV.2 Fermionic particle modes
[…]
@@
$\gamma^{\theta}(\Theta,\ell)$ $=$ $\frac{1}{\sqrt{{C^{(\Theta,\ell)}}}}\left(\begin{array}[]{cc}\bf{0}&\sigma_{1}\\ \sigma_{1}&\bf{0}\end{array}\right),\;\;\gamma^{\varphi}(\Theta,\ell)=\frac{1}{{\sqrt{D^{(\Theta,\ell)}}}}\left(\begin{array}[]{cc}\bf{0}&\sigma_{2}\\ \sigma_{2}&\bf{0}\end{array}\right).$
@@
To represent a Dirac field whose spin is oriented upward along the positive rr–direction, the following ansatz is employed vanzo2011tunnelling :
@@
The analysis centers on the spin–up $(+)$ configuration, corresponding to alignment along the positive rr–axis. The spin-down ()(-) case, which represents orientation in the opposite direction, can be handled through an analogous procedure. Inserting the spinor ansatz from Eq. (61) into the Dirac equation yields the following set of expressions, in line with the method outlined by Vanzo et al. vanzo2011tunnelling
@@
$\begin{split}n_{\psi}(\Theta,\ell)&={\frac{{\Gamma}_{\psi}(\Theta,\ell)}{1+{\Gamma}_{\psi}(\Theta,\ell)}}\\ &=\,\,\frac{1}{\exp\left(\frac{128\pi(\ell-1)^{5}M^{3}\omega}{16(\ell-1)^{3}M^{2}-\Theta^{2}\left(9\ell^{2}+5\right)}\right)+1}\\ &{\approx\,\,\frac{16\pi\ell M\omega e^{8\pi M\omega}}{\left(e^{8\pi M\omega}+1\right)^{2}}+\frac{1}{e^{8\pi M\omega}+1}}\\ &{+\left[\frac{5\pi\ell\omega e^{8\pi M\omega}\left(16\pi M\omega e^{8\pi M\omega}-16\pi M\omega+e^{8\pi M\omega}+1\right)}{2M\left(e^{8\pi M\omega}+1\right)^{3}}+\frac{5\pi\omega e^{8\pi M\omega}}{2M\left(e^{8\pi M\omega}+1\right)^{2}}\right]\Theta^{2}}.\end{split}$ (77)
Fig. 7 displays the fermionic particle creation density, $n_{\psi}$222 Here, it is worth noting that although the expansion in the third line of Eq. (77) is presented up to second order in $\Theta$ and first order in $\ell$, the plots in Fig. 7 were generated using the full expression. This choice was made to render the effects of non–commutativity and Lorentz violation more pronounced and easier to distinguish in the figures. . As observed in the bosonic case, both the non–commutative parameter $\Theta$ and the Lorentz–violating parameter $\ell$ enhance the overall particle production rate, leading to a noticeable increase in the density’s amplitude therefore.
@@ V Evaporation and Emission rate
$\frac{\mathrm{d}M}{\mathrm{d}t}=-\alpha a{\tilde{\sigma}}T^{4},$ (78)
with $\tilde{\sigma}$} corresponds to the effective radiating (cross–section) area, $a$ is the radiation constant, $\alpha$ accounts for the greybody factor, and $T$ denotes the Hawking temperature. It is important to highlight that the emitted radiation predominantly consists of massless particles, such as photons and neutrinos hiscock1990evolution ; page1976particle . The cross–sectional area $\sigma$ is approximated by $\pi R_{sh}^{2}$, where $R_{sh}$ denotes the shadow radius, serving as the effective emission radius in the high–frequency (geometric optics) limit—often referred to as ${\tilde{\sigma}}_{\text{lim}}$. Under these conditions, the greybody factor simplifies to $\alpha\to 1$, as noted in liang2025einstein . Moreover, after some algebraic manipulations, we can write $\sigma_{lim}$ as follows:
${\tilde{\sigma}}_{lim}\approx\,27\pi M^{2}-81\pi\ell M^{2}-\frac{3\pi\Theta^{2}}{4}+\frac{3}{4}\pi\Theta^{2}\ell,$ (79)
@@
[…]
where it is expanded up to second order in $\Theta$ and first order in $\ell$ so that the expression takes a simplified form; here, we have set $\xi=a\alpha$. The next step involves evaluating the following integral:
@@
Figure 9: The evaporation lifetime at the its final stage $M_{f}\to 0$ for different values of initial mass $M_{i}$ by considering different values of $\Theta$ (up panel) and different values of $\ell$ (bottom panel).
@@
Quantum effects manifest prominently in the vicinity of a black hole’s event horizon, where transient particle–antiparticle pairs emerge due to vacuum fluctuations. Occasionally, one of these particles—carrying positive energy—may traverse the potential barrier and escape via quantum tunneling. This escape mechanism leads to a net energy loss from the black hole, contributing to its gradual evaporation, a process commonly known as Hawking radiation. For an observer located far from the black hole, this evaporation is indirectly associated with the absorption profile at high energies. In such regimes, the absorption cross–section tends toward a limiting value, denoted by ${\tilde{\sigma}}_{\text{lim}}$. As reported in Refs. decanini2011universality ; papnoi2022rotating , this behavior governs the spectral energy emission, which can be formulated as follows:
$\frac{{{\mathrm{d}^{2}}E}}{{\mathrm{d}\omega\mathrm{d}t}}=\frac{{2{\pi^{2}}{\tilde{\sigma}}_{lim}}}{{{e^{\frac{\omega}{T^{(\Theta,\ell)}}}}-1}}{\omega^{3}},$ (84)
@@
$\frac{\mathrm{d}^{2}N}{\mathrm{d}\omega\mathrm{d}t}=\frac{2\pi^{2}\,{\tilde{\sigma}}_{lim}\,\omega^{2}}{e^{\,\omega/T^{(\Theta,\ell)}}-1}=\frac{\pi^{3}\omega^{2}\left(\Theta^{2}(\ell+2)+72(3\ell-2)M^{2}\right)^{2}}{384M^{2}\left(e^{-\frac{128\pi(\ell-1)^{3}M^{3}\omega}{3\Theta^{2}+16(\ell-1)^{2}(\ell+1)M^{2}}}-1\right)}.$ (86)
@@ VI Massless Scalar Field in a Non-Commutative Background
Substituting the Eqs. (106)-(107) into the Eq. (105) and considering the new redefined radial function $\Psi_{lm}(r^{*})$ in the Klein–Gordon equation, a one–dimensional Schrödinger–like wave equation will be obtained as
@@
$\left.\left.+768r^{3}\left(\frac{f^{4}(r)}{2\sqrt{f(r)}}-\frac{3f^{\prime\prime}(r)^{2}}{4f(r)^{3/2}}-\frac{15f^{\prime}(r)^{4}}{16f(r)^{7/2}}-\frac{f^{3}(r)f^{\prime}(r)}{f(r)^{3/2}}+\frac{9f^{\prime}(r)^{2}f^{\prime\prime}(r)}{4f(r)^{5/2}}\right)\right)\right)\Bigg{]}.$ (111)
@@
where the primes denote derivatives with respect to the radial coordinate $r$. It is worth noting that the effective potential $V_{\text{eff}}(r)$ is well defined only for $l>0$. Accordingly, since Eq. (111) was obtained for $l=1$, taking the limit $\Theta\to 0$ clearly recovers the standard effective potential for a scalar field in the Kalb–Ramond black hole background with $l=1$ araujo2024exploring , as expected.
@@ VII Quasinormal modes
Additional perspective is got by evaluating the dependence of the deviations on the overtone number $n$, as depicted in Figs. 1920. The real parts of the frequencies, which correspond to oscillation rates, increase with $\Theta$ more significantly for higher overtones. The imaginary parts—which govern damping—also show a rising trend with $\Theta$, but the most prominent changes occur for the fundamental mode $(n=0$). This suggests that lower overtones are more sensitive to non–commutative corrections in terms of energy dissipation.
@@ VIII Time–domain solution
Here, the evolution begins by assigning the scalar field a vanishing value along the initial surface $v=v_{0}$, that is, $\psi(u,v_{0})=0$. This choice simplifies implementation without restricting the general applicability of the method. The numerical integration proceeds by holding $u$ constant while incrementally advancing in the $v$ direction, making use of the null boundary data. For the sake of simplicity, the scalar perturbation is treated in a background where $M=1$. The initial waveform is introduced as a Gaussian pulse centered at $v=0$, with a fixed width $\sigma=1$ and zero amplitude at the origin. The computational grid is defined over the square domain $0,1000]\times[0,1000]$ in the $(u,v)$ plane, using a uniform spacing of $h=0.1$.
The progression of the scalar field $\psi$ over time is illustrated in Fig. 21, where distinct values of the non–commutative parameter $\Theta$—ranging from $0.1$ to 0.40.4—are examined under the constraint $\ell=0.1$. The behavior is evaluated across different angular modes: $l=1$ (upper left), $l=2$ (upper right), and $l=3$ (bottom), with the azimuthal number fixed at $m=+1$. The evolution profiles indicate that the field tends to stabilize as time increases, eventually flattening out to small values.
@@
Figure 21: The time–dependent behavior of the field $\psi$ is illustrated for different angular indices, with panels corresponding to $l=1$ (top left), $l=2$ (top right), and $l=3$ (bottom), under the fixed choice of $m=+1$. The parameter $\Theta$ is varied across the set 0.1,0.2,0.3,0.4{0.1,0.2,0.3,0.4} to examine its influence on the wave evolution, while the value of $\ell$ is kept unchanged at $0.1$ throughout the analysis.
@@ IX Estimated constraints from the solar system tests
Moreover, the expressions for the conserved quantities associated with the system—namely, the energy $E$ and the angular momentum LL—can be written as follows
@@ IX.1 Perihelion precession of Mercury
$u_{0}=1+e\cos(\varphi)\,$, (124)
where $e$ is the eccentricity of the orbit. Here, by inserting the expanded form of $u$ into Eq. (123), one arrives at the following differential equation structure:
@@
When considering the correction associated with the $\epsilon$ parameter in Eq. (125), its influence must vanish to preserve the perturbative structure, yielding new corrections. Among the various sinusoidal terms that arise—most of which average to zero over a full orbit—only one term, $\frac{1}{2}e\varphi\sin(\varphi)$, contributes meaningfully by producing a secular variation. After accounting for the Newtonian background, the relativistic correction, and this new contribution, and by reintroducing the original theoretical parameters, the resulting expression for $u(\varphi)$ becomes:
@@
The angular momentum $L$ of a planetary orbit can be expressed in terms of the semi-major axis $a$ and eccentricity $e$ through the relation $L^{2}=Ma(1-e^{2})$, while the specific orbital energy per unit mass is given by E=M/(2a)E=-M/(2a) Goldstein:2002 . In natural units, the physical parameters for Mercury’s orbit and the Sun are set as follows: the solar mass is $M=M_{\odot}=9.138\times 10^{37}$, the semi–major axis of Mercury’s orbit is $a=3.583\times 10^{45}$, and its eccentricity is $e=0.2056$. Using these values, one obtains $L=5.600\times 10^{41}$, confirming the validity of a perturbative approach due to the smallness of $M^{2}/L^{2}$ contributions. The squared energy, calculated as $E^{2}=1.627\times 10^{-16}$, is small enough that energy–dependent corrections can safely be disregarded.
With the parameter $\epsilon$ defined by $\epsilon=-\ell-\frac{\Theta^{2}}{2L^{2}}(1-E^{2})$, one can derive independent constraints on $\ell$ and $\Theta$. Setting $\Theta=0$ isolates the Kalb–Ramond contribution, leading to the constraint $1.817\times 10^{-11}\leq\ell\leq 3.634\times 10^{-12}$. Alternatively, taking $\ell=0$ allows for bounding the non–commutative parameter, resulting in the range $2976.57\,\text{m}^{2}\leq\Theta^{2}\leq 595.315\,\text{m}^{2}$. For completeness, both positive and negative branches of $\Theta^{2}$ are retained in the analysis.
@@ IX.2 Deflection of light
In this context, the ratio $L/E$ corresponds to the impact parameter $b$. To isolate the effects ascribed to $\Theta$, we first set $\ell=0$. Notably, the resulting corrections involving the parameter $\Theta$ are universal—they appear independently of the central mass MM—and thus modify the Newtonian trajectory from the outset. By appropriately redefining the mass term, the leading–order contributions can be expressed as follows:
@@
General Relativity predicts a deflection angle of $\delta_{\text{GR}}=4M/b=1.7516687^{\prime\prime}$. Observationally, the measured deflection is expressed as $\delta_{\text{Exp}}=\frac{1}{2}(1+\gamma)\times 1.7516687^{\prime\prime}$, with $\gamma=0.99992\pm 0.00012$ dsasdas . To examine the influence of $\Theta$, the dimensionless factor $1-\Theta^{2}/(4b^{2})$ is compared with the empirical ratio $(1+\gamma)/2$. This comparison yields the following constraint: $3.872\times 10^{13}\,\text{m}^{2}\leq\Theta^{2}\leq 1.936\times 10^{14}\,\text{m}^{2}$.
@@
By assuming the regime of small angular deviations, $\varphi\ll 1$, and following the same reasoning outlined previously, the resulting deflection angle in the presence of the Kalb–Ramond background becomes $\delta_{\ell}=\frac{4M}{b}(1-\frac{3}{2}\ell)$. This expression allows us to infer the constraint on $\ell$, yielding the range $1.333\times 10^{-5}\leq\ell\leq 6.667\times 10^{-5}$.